# RASTP: Representation-Aware Semantic Token Pruning for Generative Recommendation with Semantic Identifiers

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
@@ 1. INTRODUCTION
Generative Recommendation (GR) has recently emerged as a powerful paradigm in industrial recommender systems, primarily due to its ability to perform end-to-end next-item prediction by directly generating item identifiers. Central to this paradigm is the adoption of Semantic Identifiers (SIDs), which represent each item as a hierarchical sequence of semantic codewords (typically three to five) (onerec; rqvae). This design enables semantically similar items to share representations and supports a compact encoding of an exponentially large item space without a massive vocabulary.
Pioneered by TIGER (tiger), which adapted transformer-based sequence modeling to predict item SIDs for recommendation, subsequent works have enhanced SID-based GR using auxiliary information or advanced architectures (other2; other3). Recent efforts have introduced standardized frameworks (grid) and industrial-scale benchmarks (forge) to advance SID-based generative recommendation.
Despite their strong performance, using SIDs incurs significant computational overhead. The representation of SIDs made of multiple tokens introduces substantially longer input sequences. This expansion leads to increased training times, which constitutes a critical bottleneck in industrial settings, where models must be retrained daily on billions of new interactions to remain up-to-date. While existing works primarily target attention computation (fuxi), KV caching (earn), or architecture (flame), direct optimization of the SIDs remains underexplored, leading to a significant waste of computational resources on meaningless calculations.
Therefore, to address the limitations of existing methods, we focus on observing semantic information and find that not all semantic features of items contribute equally to predictions, as users typically only attend to a subset of them. This phenomenon is reflected in the tokens. After processing through Transformer layers, certain tokens become redundant because their content has already been represented by more informative tokens. Motivated by these observations and related works (H2O; VATP), we propose RASTP (Representation-Aware Semantic Token Pruning), a efficient strategy that dynamically identifies and retains more informative SIDs. RASTP evaluates token importance based on attention weights and intermediate representations from Transformer layer, enabling the model to dynamically focus on the most discriminative SIDs.
@@ 2.1. Semantic ID Tokenization
This dense embedding is then discretized into a hierarchical sequence of SID via a quantization tokenizer. We employ RQ-Kmeans (onerec), a hierarchical vector quantization method that maps $\mathbf{h}_{i}$ to an $L$-level SID sequence:
@@ 3.1. Experimental Settings
We conduct our evaluation on the 5-core filtered Amazon datasets for Beauty, Sports, and Toys. For each user, we reserve the last interaction as the test sample, the second-to-last for validation, and all earlier interactions for training. Item textual features consist of Title, Categories, Description, and Price.
@@
For SID tokenization, item textual features are encoded via mean pooling over the final hidden states of Flan-T5-XL. We use RK-Means with three codebooks by default, as the quantization method does not affect our core findings. The generative recommendation model employs an encoder-decoder architecture with 8 Transformer layers (4 encoder + 4 decoder). All other hyperparameters are configured as detailed in Table 1.
@@
Generative
Recommendation
(T5 Encoder-Decoder)
GPU NVIDIA A100-PCIE-40GB
@@
3.2. Performance of Generative Recommendation with RASTP (RQ1)
@@ 3.3. Different Pruning Strategies (RQ2)
To investigate how different token pruning strategies affect recommendation performance, we compare RASTP against two baselines: (1) simple pooling methods, specifically max and average pooling, and (2) an information-aware selection strategy based on the $\ell_{2}$ norm of token representations. Results in Figure 2 show that pooling methods cause a substantial performance drop due to excessive loss of fine-grained semantic information. Although $\ell_{2}$-norm-based pruning performs better, it still lags behind RASTP. In contrast, RASTP consistently maintains strong performance across all evaluation metrics, demonstrating that its representation-aware token selection effectively preserves semantic fidelity while enabling sequence compression—thereby sustaining both model expressiveness and recommendation quality.
@@ 3.4. Performance or Efficiency (RQ3)
Generative recommendation models typically employ multi-layer Transformer architectures, in which the timing of token pruning critically affects the trade-off between efficiency and performance. Pruning too early yields greater computational savings but risks removing informative tokens before their contextual representations are adequately refined.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792846isbn: 979-8-4007-2307-0/2026/04isbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
@@ 1. INTRODUCTION
Generative Recommendation (GR) has recently emerged as a powerful paradigm in industrial recommender systems. Unlike prior methods that merely used large models to augment traditional pipelines (Yu et al., 2025; Lv et al., 2025), GR performs end-to-end next-item prediction by directly generating item identifiers. Central to this paradigm is the adoption of Semantic Identifiers (SIDs), which represent each item as a hierarchical sequence of semantic codewords (Zhou et al., 2025; Lee et al., 2022). This design enables semantically similar items to share representations and supports a compact encoding of an exponentially large item space without a massive vocabulary.
Pioneered by TIGER (Rajput et al., 2023), which adapted transformer-based sequence modeling to predict item SIDs for recommendation, subsequent works have enhanced SID-based GR using auxiliary information or advanced architectures (Yang et al., 2025b; Tan et al., 2024). Recent efforts have introduced standardized frameworks (Ju et al., 2025) and industrial-scale benchmarks (Fu et al., 2025) to advance SID-based generative recommendation.
Despite their strong performance, using SIDs incurs significant computational overhead. The representation of SIDs made of multiple tokens introduces substantially longer input sequences. This expansion leads to increased training times, which constitutes a critical bottleneck in industrial settings, where models must be retrained daily on billions of new interactions to remain up-to-date. While existing works primarily target attention computation (Ye et al., 2025), KV caching (Yang et al., 2025a), direct optimization of the SIDs remains underexplored, leading to a significant waste of computational resources on meaningless calculations.
Therefore, to address the limitations of existing methods, we focus on observing semantic information and find that not all semantic features of items contribute equally to predictions, as users typically only attend to a subset of them. This phenomenon is reflected in the tokens. After processing through Transformer layers, certain tokens become redundant because their content has already been represented by more informative tokens. Motivated by these observations and related works (Zhang et al., 2023; Guo et al., 2024), we propose RASTP (Representation-Aware Semantic Token Pruning), a efficient strategy that dynamically identifies and retains more informative SIDs. RASTP evaluates token importance based on attention weights and intermediate representations from Transformer layer, enabling the model to dynamically focus on the most discriminative SIDs.
@@ 2.1. Semantic ID Tokenization
This dense embedding is then discretized into a hierarchical sequence of SID via a quantization tokenizer. We employ RQ-Kmeans (Zhou et al., 2025), a hierarchical vector quantization method that maps $\mathbf{h}_{i}$ to an $L$-level SID sequence:
@@ 3.1. Experimental Settings
We conduct our evaluation on the 5-core filtered Amazon datasets for Beauty, Sports, and Toys. For each user, we reserve the last interaction as the test sample, the second-to-last for validation, and all earlier interactions for training.
@@
For SID tokenization, item textual features are encoded by Flan-T5-XL. We use RK-Means with three codebooks by default, as the quantization method does not affect our core findings. The generative recommendation model employs an encoder-decoder architecture with 8 Transformer layers (4 encoder + 4 decoder). All other hyperparameters are configured as detailed in Table 1.
@@
Generative Recommendation (T5 Encoder-Decoder) GPU NVIDIA A100-PCIE-40GB
@@
3.2. Performance with RASTP (RQ1)
@@ 3.3. Different Pruning Strategies (RQ2)
To investigate how different token pruning strategies affect recommendation performance, we compare RASTP against two baselines: (1) simple pooling methods, specifically max and average pooling, and (2) an information-aware selection strategy based on the $\ell_{2}$ norm of token representations. Results in Figure 2 show that pooling methods cause a substantial performance drop due to excessive loss of fine-grained semantic information. Although $\ell_{2}$-norm-based pruning performs better, it still lags behind RASTP. In contrast, RASTP consistently maintains strong performance across all evaluation metrics, demonstrating that its representation-aware token selection effectively preserves semantic fidelity while enabling sequence compression.
@@ 3.4. Performance or Efficiency (RQ3)
Generative recommendation models typically employ multi-layer Transformer architectures, in which the timing of token pruning critically affects the trade-off between efficiency and performance.
@@ 4. CONCLUSION
5. ACKNOWLEDGMENTS
This work is supported by the 2030 National Science and Technology Major Project (2022ZD0119100).